# Density Estimation via Measure Transport: Outlook for Applications in the Biological Sciences

### v2

@@ 5.2.2 Adaptive Transport Maps as a Tool for Scientific Discovery
    $\pi_{k}$: (unknown) target probability density for Class $k\in\{0,\ldots,K-1\}$
    $Z_{k}=\{z\in\mathbb{R}^{m}\,:\,z\sim{\nu_{\pi_{k}}}\}$: set of all samples for class $k\in\{0,1,\ldots,K-1\}$
@@
    $\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}\,$: $m\times m$ lower triangular matrix storing frequency counts from sparsity
        patterns for $N$ transport maps computed for class $k$ and permutation $\sigma_{{p}}(\mathcal{{Y}})$
@@
1:Choose (at random) a permutation $\sigma_{{p}}\in\mathcal{S}_{m}$ on $m$ elements. $\triangleright$ see Table 2
@@
11:         Compute an adaptive transport map approximation $T_{k}^{(N,r)}$ from the
         reference $\rho$ to the target density ${\pi_{k}}$ using only $\sigma\left(Z_{k}\right)^{\mathrm{tr}}$ as a training set.
@@
          $\nu_{\pi_{k}}\left(\sigma\left(Z_{k}\right)^{\mathrm{tr}}\right)=\nu_{\rho}\left(\left(T_{k}^{(N,r)}\right)^{-1}\!\left(\sigma\left(Z_{k}\right)^{\mathrm{tr}}\right)\right)$. (cf. equation (2))
@@
Algorithm 2 Gathering statistics from transport maps sparsity patterns
@@
Namely, like in Algorithm 1 before, we work in Algorithm 2 with subsets (as training sets) extracted from the set of data samples after random permutations of the ordered set of GSE43151 data samples (see lines 810 of Algorithm 2). First, we fix an ordering $\sigma_{{p}}(\mathcal{{Y}})=\left(\mathcal{{Y}}_{\sigma_{{p}}^{-1}(j)}\right)_{j=1}^{6}$ on the set $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ of random variables (i.e., we fix an ordering on the arguments of the transport maps to be computed), where $\sigma_{{p}}\in\mathcal{S}_{6}$ – the permutation group on six ($m=6$) elements. Recall that we consider only triangular maps throughout (see equations 1017). We restrict our attention to only two ($K=2$) classes: the zero-dose ($k=0$) and low-dose ($k=1$) radiation classes. For each class $k\in\{0,1\}$ and a given number $N$, we compute a series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$ of transport maps $T_{k}^{(N,r)}\!:\mathbb{R}^{6}\rightarrow\mathbb{R}^{6}$, $r=1,\ldots,N$, rendering approximations $\pi_{k}^{(N,r)}$ to the target density $\pi_{k}$ on $\mathbb{R}^{6}$ describing the joint density distribution of the six random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ in question. More precisely, for each class $k\in\{0,1\}$ we performed five independent computations of series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$ of transport maps with a number of maps (runs) $N=10,20,50,100,200$ in each, respectively, and with the same fixed order $\sigma_{{p}}(\mathcal{{Y}})$ of the random variables.
@@
We observed that for each class $k\in\{0,1\}$ (and more so with an increasing $N$) the elements in the series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$ cluster around a map $\overline{T}_{k}$ the components $\{(\overline{T}_{k}^{-1})_{i}\}_{i=1}^{6}$ of whose inverse show a particular persistent and dominant pattern of dependence on its arguments (active variables), referred to previously as the sparsity pattern (see Figure 3 for the definition). In order to exhibit this pattern quantitatively, for each of the zero-dose and low-dose radiation classes, we recorded the frequency with which each random variable in $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ appeared as an active variable in the components $\{(\left(T_{k}^{(N,r)}\right)^{-1})_{i}\}_{i=1}^{6}$, separately within each of the five independently computed series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$ of adaptive transport maps, for $N=10,20,50,100,200$ as explained above. This corresponds to lines 1417 in Algorithm 2. Using these frequency counts we then constructed an “aggregated” sparsity pattern for each of the five series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$, $N=10,20,50,100,200$, for each one of the two classes $k\in\{0,1\}$. Results are shown in Figures 10 and 11, respectively, for the zero-dose and low-dose radiation classes. To simplify notation, whenever there is no danger of confusion, in what follows we drop the subscript $k$ and superscript $(N,r)$ in the transport map notation $T_{k}^{(N,r)}$ introduced above and in Algorithm 2.
@@
Figure 10: Aggregated sparsity patterns from adaptive transport map approximations to data distributions for different subsets of the zero-dose radiation data samples. The permutation on the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables for the computational experiments summarized in Subfigures 10a10f is the identity permutation $\sigma_{{1}}(\mathcal{{Y}})=(\mathcal{{Y}}_{1}\ \mathcal{{Y}}_{2}\ \mathcal{{Y}}_{3}\ \mathcal{{Y}}_{4}\ \mathcal{{Y}}_{5}\ \mathcal{{Y}}_{6})$ (see Table 2 from Appendix B). Subsets of samples consisting of half the number of total samples (which for the zero-dose radiation class is 18) were generated from random permutations of the GSE43151 data set sample ordering. In Subfigures 10a, 10b, 10c, 10d, and 10e are represented the sparsity patterns corresponding to, respectively, 10, 20, 50, 100, and 200 such sample subsets. By aggregating results for increasing numbers of sample subsets, patterns of high-valued frequency counts are seen to emerge and persist. The groups of runs represented in Subfigures 10a10e were independent, so low-valued frequency counts in the sparsity patterns need not persist as we move from Subfigure 10a to Subfigure 10e. This was the case for grid point (5,6) in Subfigures 10d and 10e, which shows a frequency count of two on the value $y_{5}$ of the random variable $\mathcal{{Y}}_{5}$ for map component $S_{6}$ in Subfigure 10d, but a frequency count of zero in Subfigure 10e. For reference, Subfigure 10f shows the frequency counts corresponding to the sparsity pattern from Subfigure 10e. In other words, the matrix shown in Subfigure 10f is the matrix $\mathrm{{B}}_{{0}}^{{(200,\sigma_{{1}})}}$ from Algorithm 2, lines 1219. To simplify notation in Subfigure 10f, we denote $S\equiv T^{-1}$ for the inverse $T^{-1}$ of a transport map $T$.
@@
Computing frequency counts for (independently computed) larger collections of transport maps, as in Figures 10a10e for the zero-dose radiation class and Figures 11a11e for the low-dose one, we observe that patterns of high-valued frequency counts (pointing to particular subsets of random variables) start to emerge and persist. We note that a larger collection of transport maps need not contain the smaller previously independently computed collection of transport maps, for, as we know from the description of Algorithm 2, the computation of each transport map uses as a training set a randomly selected subset of the set of all available samples for the given class – that is indeed the case in Figures 10 and 11.
In order to verify robustness of the results with respect to changes in the ordering of the random variables, the computational experiments described via Figures 1011 were repeated using $20$ different permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ of the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables. The different permutations used are listed in Table 2, appearing in Appendix B. Details about the experiments for each of the permutations are provided in Appendix B. Here we note that the persistence of patterns of high-valued frequency counts resulting from the computational experiments, as described in Figures 1011, and which correrspond to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$ case, was also observed in the case of each of the 19 permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=2}^{20}$.
We thus take the results from the case of $N=200$ runs (cf. Figures 10e and 11e) for each of the $20$ permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$, and map them to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$, so that we can report the aggregated results for these $4,\!000$ runs. This is summarized in Figures 1213. The resulting high-valued frequency counts in the sparsity patterns are suggestive of gene relationships encoded (or hidden) in the data which, as we have already mentioned, may aid in hypothesis generation for subsequent validation.
@@
Figure 13: Subfigures 13a13b: Dependencies among the genes for the zero-dose radiation class, as implied by the sparsity pattern from Figure 12a and the associated frequency counts from Figure 12b. Subfigures 13c13d: Dependencies among the genes for the low-dose radiation class, as implied by the sparsity pattern from Figure 12c and the associated frequency counts from Figure 12d. Information about the subset of six genes selected for the computational experiments, including the gene IDs, appears in Table 1 from Appendix B. The entries in the tables shown in Subfigures 13b and 13d above contain, respectively, the frequency counts from Figures 12b and 12d as a fraction of a total number of $4,\!000$ runs, and mapped to the gene IDs. These fractional values and the (user defined) thresholds from the legend at the top right corner of Subfigure 13b are used in the above figures to represent a “strength” of the relationship between genes, as implied by the adaptive transport maps sparsity patterns from Figures 12a and 12c. A natural question is whether the above reveal biological information, such as relationships among genes with no radiation exposure that change (appear, disappear) or persist upon exposure to low-dose radiation.
@@ 6 Discussion
As mentioned in Section 1, one among several advantages of measure transport methods is that they allow for a unified framework for processing and analysis of data distributed according to a wide range of probability measures. This powerful property is one of the reasons why measure transport techniques are active areas of research up to the present day, despite the fact that the subject itself has a long history. A framework for probability density estimation supported by measure transport techniques clearly offers benefits for many applications, including those in biology. For instance, recent work by [7] has shown that improved interpretation of cancer transcriptomic data can result by accounting for differences in the distribution densities of gene expressions. This is just one example where a framework for probability distribution density estimation, grounded on measure transport methods, could have significant impact.
@@
Specific to the present manuscript, one of the ultimate goals when studying the GSE43151 data set is to understand the global dependence structure between the genes represented there through their expression data interpreted as a set of values of a set of random variables with a joint $d$-variate probability density $\Pi_{k}$ on $\mathbb{R}^{d}$, where $d$ denotes the number of genes in the data set, and $k\in\{0,1,\ldots,K-1\}$ labels the different radiation level classes. The GSE43151 data set we used (available from [21]) contains samples for $d=10,\!874$ genes. The approach and methods we presented here (which are inclusive of, although not exclusively for, scarce data scenarios) are very general and to our knowledge have not been pursued in the biology domain. (We note, however, that application of other measure transport methods in the biology domain appear in the literature – for example, see [43, 6, 13] for neural network and optimal transport applications.) Combined with present day computational capabilities, such an approach renders the task of achieving goals as the above feasible.
@@
For instance, within this framework, the computations and results from Section 5.2.2 can be viewed as follows. Choosing a subset of $m$ genes (where $m=6$ genes were considered in Section 5.2.2) we essentially chose a subspace $\mathbb{R}^{m}\subset\mathbb{R}^{d}$ and, by computing triangular transport maps $T_{k}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, estimated target probability densities $\pi_{k}$ on $\mathbb{R}^{m}$ which can be thought of as approximations of $m$-variate marginals of the (unknown) global $d$-variate probability densities $\Pi_{k}$. In other words, we chose a (known) projection $P:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ and estimated the marginals $\nu_{\pi_{k}}=P_{\sharp}{\nu_{\Pi_{k}}}$ (transports of $\nu_{\Pi_{k}}$ along this projection) on $\mathbb{R}^{m}$ by constructing triangular transport maps satisfying $\nu_{\pi_{k}}=(T_{k})_{\sharp}\nu_{\rho}$, where the reference probability density $\rho$ on $\mathbb{R}^{m}$ was chosen to be an $m$-variate Gaussian.
@@
The dependence structure between genes in this small subset (as implied by the computed estimates to the probability distribution densities) was then uncovered by analyzing the sparsity patterns exhibited by the computed triangular transport maps. In general, assuming that $\mathbb{R}^{d}$ is equipped with a reference Gaussian measure $\nu_{\Gamma}$, one may address the question of computing global triangular transport maps $\mathcal{T}_{k}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, satisfying $\nu_{\Pi_{k}}=(\mathcal{T}_{k})_{\sharp}\nu_{\Gamma}$ and estimating the (unknown) global probability density $\Pi_{k}$, where $\Gamma$ is a $d$-variate Gaussian reference probability density. It will be interesting to understand the sparsity patterns of the global maps $\mathcal{T}_{k}$, as their structure will reveal dependencies and strength of correlations between different genes represented in the entire data set.
@@
Note that such an insight should reveal known and unknown information about pathways as well, since pathway information should be encoded into the dependency structure of the global target probability densities $\Pi_{k}$ and therefore in the sparsity patterns of the global triangular transport maps $\mathcal{T}_{k}$. As already mentioned, in Section 5.2 we do not assume a priori any particular structure of the gene network. For example, graphs that represent (portions of) a gene network are not acyclic, in general. Thus, our method and analysis from Section 5.2.2 is not restricted a priori to directed acyclic graphs assumed to represent gene regulatory networks.
@@
The results of adopting this general approach are already seen in Figure 13, where emerging complex relations can be observed in the low-dose radiation class within the small subset of six genes (i.e., random variables) that we studied in detail. Whether to ignore or not the impact of possible feedback loops in the gene network, as well as the overall interpretation of the relationships among genes uncovered from the transport maps sparsity patterns, will depend on an external expert knowledge and judgment. Extracting local information by estimating the triangular transport maps ${T}_{k}$ and target densities $\pi_{k}$ on a subset of genes and/or using an already established knowledge about the relations between genes within a subset (for example, a pathway known with certainty, which will predetermine the sparsity pattern of local transport maps and the dependency structure of local target probability densities) can help in computing the global transport maps $\mathcal{T}_{k}$ and estimating target densities $\Pi_{k}$. Namely, this prior knowledge can be incorporated into the original mathematical problem (i.e., first equation in the system (6) below) as added constraints:
$\nu_{\Pi_{k}}$ $=$ $(\mathcal{T}_{k})_{\sharp}\nu_{\Gamma}$
$(T_{k})_{\sharp}\left(P_{\sharp}\nu_{\Gamma}\right)$ $=$ $P_{\sharp}\left((\mathcal{T}_{k})_{\sharp}\nu_{\Gamma}\right)$
@@
where the local map ${T}_{k}$ has been computed already and the second equation in the system (6) above (i.e., constraint for the global triangular transport map $\mathcal{T}_{k}$) follows from the commutativity condition $T_{k}\circ{P}=P\circ\mathcal{T}_{k}$.
@@
Furthermore, the results depicted in Figure 13 suggest a certain “dynamics” on the space of transport maps (reflected in the changes to their sparsity patterns), as the level of radiation for class $k\in\{0,1,\ldots,K-1\}$ changes, revealing also changes in correlations between genes in the subset of genes under consideration. Moreover, as described in [30], one could gather gene expression data within a fixed class $k\in\{0,1,\ldots,K-1\}$ at different instances of time $t=t_{1},t_{2},\ldots,t_{n}$ after an irradiation with a certain dose (labeled by the class $k$) has occurred. In other words, one observes the time dynamics on the space of random variables $\mathcal{{Y}}(t)=\left(\mathcal{{Y}}_{1}(t),\mathcal{{Y}}_{2}(t),\dots,\mathcal{{Y}}_{d}(t)\right)\in\mathbb{R}^{d}$, which implies time dynamics on the space of target probability measures $\nu_{\Pi_{k}}(t)$ (or equivalently, on the space of target probability densities ${\Pi_{k}}(t)$), which itself induces time dynamics on the space of triangular transport maps $\mathcal{T}_{k}(t)$ estimating the latter, all within the class $k$.

### v4

@@ 5.2.2 Adaptive Transport Maps as a Tool for Scientific Discovery
    $\psi_{k}$: (unknown) target probability density for Class $k\in\{0,\ldots,K-1\}$
    $Z_{k}=\{z\in\mathbb{R}^{m}\,:\,z\sim{\nu_{\psi_{k}}}\}$: set of all samples for class $k\in\{0,1,\ldots,K-1\}$
@@
    $\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}\,$: $m\times m$ lower triangular matrix storing frequency counts from sparsity patterns for
        the $N$ adaptive transport maps learned for class $k$ and permutation $\sigma_{{p}}(\mathcal{{Y}})$
@@
1:Choose a permutation $\sigma_{{p}}\in\mathcal{S}_{m}$ on $m$ elements. $\triangleright$ see Table 2
@@
11:         Learn an adaptive transport map approximation $T_{k}^{(N,r)}$ from the
         reference $\rho$ to the target density ${\psi_{k}}$ using only $\sigma\left(Z_{k}\right)^{\mathrm{tr}}$ as a training set.
@@
          $\nu_{\psi_{k}}\left(\sigma\left(Z_{k}\right)^{\mathrm{tr}}\right)=\nu_{\rho}\left(\left(T_{k}^{(N,r)}\right)^{-1}\!\left(\sigma\left(Z_{k}\right)^{\mathrm{tr}}\right)\right)$. (cf. equation (5))
@@
Algorithm 2 Gathering statistics from adaptive transport maps sparsity patterns
Namely, like in Algorithm 1 before, we work in Algorithm 2 with subsets (as training sets) extracted from the set of data samples after random permutations of the ordered set of GSE43151 data samples (lines 810 in Algorithm 2). First, we select at random an ordering
$\sigma_{{p}}(\mathcal{{Y}})$ $=$ $\left(\mathcal{{Y}}_{\sigma_{{p}}^{-1}(j)}\right)_{j=1}^{6}$ (27)
on the set $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ of random variables and then fix this ordering (i.e., we fix an ordering on the arguments of the transport maps to be learned – this corresponds to lines 12 in Algorithm 2), where in (27) above we have $\sigma_{{p}}\in\mathcal{S}_{6}$ – the permutation group on six ($m=6$) elements. Recall that we consider only triangular maps throughout (see equations (15)–(22)). Also, for the purpose of illustration, we restrict our attention to only two ($K=2$) classes: the zero-dose ($k=0$) and low-dose ($k=1$) radiation classes.
@@
For each class $k\in\{0,1\}$ and a given number $N$, (see loop starting in line 3 of Algorithm 2) we learn a series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$ of adaptive transport maps $T_{k}^{(N,r)}\!\!\!:\!\mathbb{R}^{6}\!\rightarrow\!\mathbb{R}^{6}$, $r=1,\ldots,N$, rendering approximations $\psi_{k}^{(N,r)}$ to the target density $\psi_{k}$ on $\mathbb{R}^{6}$ describing the joint density distribution of the six random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ in question. More precisely, for each class $k\in\{0,1\}$ we performed five independent computations of series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$ of transport maps by executing Algorithm 2, independently, with $N\in\{10,20,50,100,200\}$ (see line 3 in Algorithm 2), and with the same fixed order $\sigma_{{p}}(\mathcal{{Y}})$ (27) of the random variables.
In doing so, we observed that for each class $k\in\{0,1\}$ (and more so with an increasing $N$) the elements in the series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$ of adaptive transport maps cluster around a map $\overline{T}_{k}$ the components $\{(\overline{T}_{k}^{-1})_{i}\}_{i=1}^{6}$ of whose inverse show a particular persistent and dominant pattern of dependence on its arguments, or active variables, which define a sparsity pattern for the inverse of the map $\overline{T}_{k}$.
@@
In order to exhibit this dominant sparsity pattern quantitatively, for each of the zero-dose and low-dose radiation classes, we recorded the frequency (line 16 in Algorithm 2) with which each random variable in $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ appeared as an active variable in the components $\{(\left(T_{k}^{(N,r)}\right)^{-1})_{i}\}_{i=1}^{6}$, separately within each of the five independently computed series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$ of adaptive transport maps, for $N=10,20,50,100,200$ as explained above. This corresponds to lines 1219 in Algorithm 2. Using these frequency counts we then constructed an “aggregated” sparsity pattern for each of the five series $\{T_{k}^{(N,r)}\}_{r=1}^{N}$, $N=10,20,50,100,200$, for each one of the two classes $k\in\{0,1\}$. The results are shown in Figures 10 and 11, respectively, for the zero-dose and low-dose radiation classes. To simplify notation in what follows, whenever there is no danger of confusion, we drop the subscript $k$ and superscript $(N,r)$ in the transport map notation $T_{k}^{(N,r)}$ introduced above and in Algorithm 2.
@@
Figure 10: Aggregated sparsity patterns from adaptive transport map approximations to data distributions for different subsets of the zero-dose radiation data samples. The permutation on the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables for the computational experiments summarized in Subfigures 10a10f is the identity permutation $\sigma_{{1}}(\mathcal{{Y}})=(\mathcal{{Y}}_{1}\ \mathcal{{Y}}_{2}\ \mathcal{{Y}}_{3}\ \mathcal{{Y}}_{4}\ \mathcal{{Y}}_{5}\ \mathcal{{Y}}_{6})$ (see Table 2 from Appendix B). Subsets of samples consisting of half the number of total samples (which for the zero-dose radiation class is 18) were generated from random permutations of the GSE43151 data set sample ordering. In Subfigures 10a, 10b, 10c, 10d, and 10e are represented the sparsity patterns corresponding to, respectively, 10, 20, 50, 100, and 200 such sample subsets. By aggregating results for increasing numbers of sample subsets, patterns of high-valued frequency counts are seen to emerge and persist. The groups of runs represented in Subfigures 10a10e were independent, so low-valued frequency counts in the depicted sparsity patterns need not persist as we move from Subfigure 10a to Subfigure 10e. This was the case for grid point (5,6) in Subfigures 10d and 10e, which shows a frequency count of two on the value $y_{5}$ of the random variable $\mathcal{{Y}}_{5}$ for map component $S_{6}$ in Subfigure 10d, but a frequency count of zero in Subfigure 10e. For reference, Subfigure 10f shows the frequency counts corresponding to the sparsity pattern from Subfigure 10e. In other words, the matrix shown in Subfigure 10f is the matrix $\mathrm{{B}}_{{0}}^{{(200,\sigma_{{1}})}}$ from Algorithm 2, lines 1219. To simplify notation in Subfigure 10f, we denote $S\equiv T^{-1}$ for the inverse $T^{-1}$ of a transport map $T$.
@@
Computing frequency counts for (independently computed) larger collections of adaptive transport maps, as in Figures 10a10e for the zero-dose radiation class and Figures 11a11e for the low-dose one, we observe that patterns of high-valued frequency counts (pointing to particular subsets of random variables) start to emerge and persist. We note that a larger collection of adaptive transport maps need not contain the smaller previously independently learned collection of adaptive transport maps, since, per lines 89 from Algorithm 2, we know that the learning of each adaptive transport map used as a training set a randomly selected subset of the set of all available samples for the given class – that is indeed the case in Figures 10 and 11.
Finally, in order to verify robustness of the results with respect to changes in the ordering of the random variables, the computational experiments described via Figures 1011 were repeated using $20$ different permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ of the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables. The different permutations used are listed in Table 2, appearing in Appendix B. Details about the experiments for each of the permutations are provided in Appendix B. Here we note that the persistence of patterns of high-valued frequency counts resulting from the computational experiments, as described in Figures 1011 for the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$ case, was also observed in the case of each of the 19 permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=2}^{20}$.
We thus take the results from the case of $N=200$ runs (cf. Figures 10e and 11e) for each of the $20$ permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$, and map them (see Algorithm 3 from Appendix B) to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$, so that we can report the aggregated results for these $4,\!000$ runs. This is summarized in Figures 1213. The resulting high-valued frequency counts in the aggregated sparsity patterns are suggestive of gene relationships encoded (or hidden) in the data which, as we have already mentioned, may aid in hypothesis generation (as was done in Figure 13) for subsequent validation.
@@
Figure 13: Subfigures 13a13b: Dependencies among the genes for the zero-dose radiation class, as implied by the sparsity pattern from Figure 12a and the associated frequency counts from Figure 12b. Subfigures 13c13d: Dependencies among the genes for the low-dose radiation class, as implied by the sparsity pattern from Figure 12c and the associated frequency counts from Figure 12d. Information about the subset of six genes selected for the computational experiments, including the gene IDs, appears in Table 1 from Appendix B. The entries in the tables shown in Subfigures 13b and 13d above contain, respectively, the frequency counts from Figures 12b and 12d as a fraction of a total number of $4,\!000$ runs, and mapped to the gene IDs. These fractional values and the (user defined) thresholds from the legend at the top right corner of Subfigure 13b are used in the above figures to represent a “strength” of the relationship between genes, as implied by the sparsity patterns from Figures 12a and 12c. A natural question is whether the above reveal biological information, such as relationships among genes with no radiation exposure (Subfigure 13a) that change (appear, disappear) or that persist upon exposure to low-dose radiation (Subfigure 13c).
@@ 6 Discussion
As mentioned in Section 1, one among several advantages of measure transport methods is that they allow for a unified framework for processing and analysis of data distributed according to a wide range of probability measures. This powerful property is one of the reasons why measure transport techniques are active areas of research up to the present day, despite the fact that the subject itself has a long history. A framework for probability density estimation supported by measure transport techniques clearly offers benefits for many applications, including those in biology. For instance, recent work [7] has shown that improved interpretation of cancer transcriptomic data can result by accounting for differences in the distribution densities of gene expressions. This is just one example where a framework for probability distribution density estimation, grounded on measure transport methods, could have significant impact.
@@
Specific to the present manuscript, one of the ultimate goals when studying the GSE43151 data set is to understand the global dependence structure between the genes represented there, through their expression data interpreted as a set of values of a set of random variables with a joint $d$-variate probability density $\Psi_{k}$ on $\mathbb{R}^{d}$, where $d$ denotes the number of genes in the data set, and $k\in\{0,1,\ldots,K-1\}$ labels the different radiation level classes. The GSE43151 data set we used (available from [21]) contains samples for $d=10,\!874$ genes. The approach and methods we presented here, which are intended to be inclusive of, although not exclusively for, scenarios where the number of samples available for training is limited in amount (order $O(10)$ samples or less), are very general and to our knowledge have not been pursued in the biology domain. (We note, however, that application of other measure transport methods in the biology domain appear in the literature – for example, see [44, 6, 13] for joint neural network and optimal transport applications.) Combined with present day computational capabilities, the methods being pursued in this paper make feasible the tasks needed to achieve goals such as that of understanding the global dependence structure among a large number of genes (or, more generally, a large number of random variables jointly distributed according to an unknown probability measure).
@@
For instance, within this framework, the computations and results from Section 5.2.2 can be viewed as follows. By choosing a subset of $m$ genes (where $m=6$ genes were considered in Section 5.2.2) we essentially chose a subspace $\mathbb{R}^{m}\subset\mathbb{R}^{d}$ and, by learning adaptive triangular transport maps $T_{k}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, estimated target probability densities $\psi_{k}$ on $\mathbb{R}^{m}$. These estimated target densities can be thought of as approximations to $m$-variate marginals of the (unknown) global $d$-variate probability densities $\Psi_{k}$. In other words, we chose a (known) projection $P:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ and estimated the marginals $\nu_{\psi_{k}}=P_{\sharp}{\nu_{\Psi_{k}}}$ (transports of $\nu_{\Psi_{k}}$ along this projection) on $\mathbb{R}^{m}$ by constructing triangular transport maps satisfying the measure transport equation $\nu_{\psi_{k}}=(T_{k})_{\sharp}\nu_{\rho}$, where the reference probability density $\rho$ on $\mathbb{R}^{m}$ was chosen to be that of the $m$-variate normal (i.e., Gaussian) measure $\mathcal{N}(0,I)$.
@@
The dependence structure between genes in this small subset (as implied by the computed estimates to the probability distribution densities) was then uncovered by analyzing the sparsity patterns exhibited by the learned adaptive triangular transport maps. In general, assuming that $\mathbb{R}^{d}$ is equipped with a reference Gaussian measure $\nu_{\Gamma}$, one may address the question of computing global triangular transport maps $\mathcal{T}_{k}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, satisfying the measure transport equation $\nu_{\Psi_{k}}=(\mathcal{T}_{k})_{\sharp}\nu_{\Gamma}$ and estimating the (unknown) global probability density $\Psi_{k}$, where $\Gamma$ is a $d$-variate Gaussian reference probability density. It will be interesting to understand the sparsity patterns of the global maps $\mathcal{T}_{k}^{-1}$, as their structure will reveal dependencies and strength of correlations between different genes represented in the entire data set.
@@
Note that such an insight should reveal known and unknown information about biological pathways as well, since pathway information should be encoded into the dependency structure of the global target probability densities $\Psi_{k}$ and therefore in the sparsity patterns of the global triangular transport maps $\mathcal{T}_{k}^{-1}$. As already mentioned, in Section 5.2 we do not assume a priori any particular structure of the gene network. For example, graphs that represent (portions of) a gene network are not necessarily acyclic, in general. Thus, our method and analysis from Section 5.2.2 was not restricted a priori to directed acyclic graphs assumed to represent gene regulatory networks.
@@
The results of adopting this general approach are already seen in Figure 13, where emerging complex relations can be observed in the low-dose radiation class within the small subset of six genes (i.e., random variables) that we studied in detail. Whether to ignore or not the impact of possible feedback loops in the gene network, as well as the overall interpretation of the relationships among genes uncovered from the adaptive transport maps sparsity patterns, will depend on an external expert knowledge and judgment. Extracting local information by estimating the triangular transport maps ${T}_{k}$ and target densities $\psi_{k}$ on a subset of genes and/or using an already established knowledge about the relations between genes within a subset (for example, a pathway known with certainty, which will predetermine the sparsity pattern of local transport maps and the dependency structure of local target probability densities) can help in learning the global transport maps $\mathcal{T}_{k}$ and thus in estimating the target densities $\Psi_{k}$. Namely, this prior knowledge can be incorporated into the original mathematical problem (i.e., first equation in the system (6) below) as added constraints (second equation in the system (6) below):
$\nu_{\Psi_{k}}$ $=$ $(\mathcal{T}_{k})_{\sharp}\nu_{\Gamma}$
$P_{\sharp}\left((\mathcal{T}_{k})_{\sharp}\nu_{\Gamma}\right)$ $=$ $(T_{k})_{\sharp}\left(P_{\sharp}\nu_{\Gamma}\right)$
@@
where the local map ${T}_{k}$ in (6) has been previously learned and the second equation in the system (6) above (i.e., constraint for the global triangular transport map $\mathcal{T}_{k}$) follows from the commutativity condition $P\circ\mathcal{T}_{k}=T_{k}\circ{P}$.
@@
Furthermore, the results depicted in Figure 13 suggest a certain “dynamics” on the space of transport maps (reflected in the changes to their sparsity patterns), as the level of radiation for class $k\in\{0,1,\ldots,K-1\}$ changes, revealing also changes in correlations between genes in the subset of genes under consideration. Moreover, as described in [30], one could gather gene expression data within a fixed class $k\in\{0,1,\ldots,K-1\}$ at different instances of time $t=t_{1},t_{2},\ldots,t_{n}$ after an irradiation with a certain dose (labeled by the class $k$) has occurred. In other words, one observes the time dynamics on the space of random variables $\mathcal{{Y}}(t)=\left(\mathcal{{Y}}_{1}(t),\mathcal{{Y}}_{2}(t),\dots,\mathcal{{Y}}_{d}(t)\right)\in\mathbb{R}^{d}$, which implies time dynamics on the space of target probability measures $\nu_{\Psi_{k}}(t)$ (or equivalently, on the space of target probability densities ${\Psi_{k}}(t)$), which itself induces time dynamics on the space of triangular transport maps $\mathcal{T}_{k}(t)$ estimating the target densities ${\Psi_{k}}(t)$, all within the class $k$.